Tasha is planning an expansion of a square flower garden in a city park. If each side of the original garden is increased by 8 meters, the new total area of the garden will be 196 square meters. Find the length of each side of the original garden.
Answer by ewatrrr(24785)

(x+8)(x+8) = 196m^2
x^2 + 16x +64 -132= 0
x^2 + 16x -132= 0 |Tossing out negative solution for unit measure
(x + 6)(x-22), x = 22, the length of each side of the original garden.